OpenStudy (unklerhaukus):

\[0.142857142\dot 857142\dot 8\]

OpenStudy (cwrw238):

nice number

OpenStudy (cwrw238):

repeating decimal - convert to a fraction?

OpenStudy (cwrw238):

- oh you haven't indicated that its repeating

OpenStudy (unklerhaukus):

ill fix that

OpenStudy (unklerhaukus):

is that right now?

OpenStudy (cwrw238):

yea - but its been a while since i did these..

OpenStudy (unklerhaukus):

multiply by one million?

OpenStudy (unklerhaukus):

\[\qquad\qquad x=0.142\dot857142\dot 8\]\[10^6x=142857.142\dot 857142\dot 8\] \[(10^6-1)x=142857\] \[x=\frac{142857}{10^6-1}=\frac{142857}{999999}\] \[\quad=\]

OpenStudy (unklerhaukus):

how can i simplify this?

OpenStudy (anonymous):

its walue of 1/7 verbly

OpenStudy (unklerhaukus):

How can i show this

OpenStudy (zepp):

>Take calculator >Enter value >Math > Frac :D

OpenStudy (unklerhaukus):

\[x=\frac{142857}{999999}=\frac{142857}{9\times111111}=\cdots\]

OpenStudy (unklerhaukus):

im dont want to have to use a calculator

OpenStudy (zepp):

111111/11 = 10101

OpenStudy (anonymous):

uper value is divisible by 9 deivide it manually

OpenStudy (unklerhaukus):

i know the 1+4+2+8+5+7=27=3x9 so the denominator but be a multiple of nine

OpenStudy (unklerhaukus):

how can i use that @zepp you mean long division @theyatin?

OpenStudy (anonymous):

i dont know what zep is thinking of but simplify whole system as much you can go for it. . .

OpenStudy (zepp):

Prime factorization :D

OpenStudy (zepp):

But you'll end up with the same thing you have above, the decimal thingy

OpenStudy (unklerhaukus):

i can not See the factors

OpenStudy (zepp):

Perform a long division with the original fraction then?

OpenStudy (unklerhaukus):